OpenStudy (anonymous):

determine the volume of the solid obtained by rotating the region bounded by x=(y-2)^2 and y=x about the line y=-1

OpenStudy (wasiqss):

well we can apply the formula \[\pi \int\limits_\int\limits_^{2dy} \]

OpenStudy (callisto):

Then you use disc method?

OpenStudy (wasiqss):

callisto first we need to take square of equation, integrate it , then put the limits

OpenStudy (callisto):

Would you mind showing the them all? (Sorry, I'm not the asker of this question though)

OpenStudy (wasiqss):

ok wait

OpenStudy (wasiqss):

squaring the eq gives, (y-2)^4 now we need to integrate which gives us \[{(y-2)^5}/5\]

OpenStudy (wasiqss):

now put the limits -1,0

OpenStudy (callisto):

Sorry, but how do you find limits to be -1, 0?

OpenStudy (wasiqss):

yeah i guess i should draw the picture

OpenStudy (wasiqss):